Night report 2021-09-15

Observers: Georgios Dimitriadis, Luke Harvey

Support: Jacco Terwel, Kate Maguire

TAT: Tom Killestein, David O'Neill, Joe Lyman 

EFOSC Calibrations:

Afternoon: EFOSC calibrations (dome flats, arcs)


Morning: EFOSC calibrations (biases)


Start: clear, humidity 20%, wind 4 m/s, seeing 1.0"

Middle: clear, humidity 20%, wind 7 m/s, seeing 1.3"

End: clear, humidity 16%, wind 8 m/s, seeing 1.0"

Comments (times in UT):

00:00 - Needed to correct the focus on the telescope, decided to retake standard afterwards.

00:54 - Switching to SOFI. Using historic focus details as the quality of the image analysis was poor. A few technical issues with SOFI that we are continuing to monitor.

03:35 - Switching back to EFOSC.